OpenStudy (anonymous):

FAN AND MEDAL! How do you solve -x^2+5x+6=0 ?

OpenStudy (anonymous):

@warpedkitten @Empty @enchanted_bubbles @RedNeckOutLaw @Teddyiswatshecallsme @yomamabf @uybuyvf @ikram002p @imqwerty @OregonDuck @phi @abb0t @sammixboo @DullJackel09 @futureboy1 @GeniousCreation @Hero @jagr2713 @KAKES1967 @Loser66 @zmudz

OpenStudy (anonymous):

plz help

OpenStudy (zmudz):

quadratic formula should do the trick, or you could factor into \(-(x-6)(x+1)\) and solve for x that way

OpenStudy (anonymous):

@zmudz i did this but its wrong -x^2+5x+6=0 (-x^2-x)+(6x+6)=0 -x(x+1)+6(x+1)=0 (-x+6)(x+1)=0 -x+6=0 -x=-6 x=6 x+1=0 x=-1

OpenStudy (zmudz):

Ah, double check your math for when you solved x=-1

OpenStudy (anonymous):

no they are both wrong i used a graphing calculator and the zeroes are x=0 and x=-1.2

OpenStudy (anonymous):

it wants me to solve this system algebraically

OpenStudy (anonymous):

Lol your answers x=6 and x=1 are both correct.

OpenStudy (anonymous):

@phi i don't know what i did wrong in solving this

OpenStudy (anonymous):

If they want you to solve it algebraically, then why did you use a graphing calculator? :D

OpenStudy (phi):

I would first multiply both sides and all terms by -1 (because I don't like -x^2) what do you get ?

OpenStudy (anonymous):

because the second part is for me to graph it to prove my answers

OpenStudy (anonymous):

x^2-5x-6=0 @phi

OpenStudy (phi):

now try factoring it

OpenStudy (anonymous):

How I wish we were allowed to use a graphing calculator when we graph something. Do you know how to plot points like this? 

OpenStudy (phi):

but it looks like you did it correctly up top x= 6 and x=-1

OpenStudy (anonymous):

(x^2-2x)-(3x-6)=0 x(x-2)-3(x-2)=0 (x-3)(x-2)=0 x=3 and x=2

OpenStudy (anonymous):

wait i messed up on the sign there

OpenStudy (anonymous):

ok I'm confused @phi

OpenStudy (anonymous):

This is if you graph it

OpenStudy (phi):

(x^2 -6x) + (x-6) = 0

OpenStudy (phi):

ok, so \[ x^2 = 5x+6 \\ x^2 -5x-6=0\\ (x-6)(x+1)= 0 \\ x= 6 ; x= -1 \] y= 5x+6 so at x= -1, y= 1 at x= 6 , y= 36 here is the graph with the first solution showing

OpenStudy (phi):

the solution is where the two curves cross (intersect) it is true when we are given a single curve (or line), we are often interested in where it crosses the x-axis (the x value is called the zero) but not in this case.

OpenStudy (anonymous):

and thats all?

OpenStudy (phi):

solving two equations is the same as asking "what point lies on both curves"

OpenStudy (phi):

there are two solutions -1,1 and (6,36)

OpenStudy (anonymous):

ok thank you!

OpenStudy (phi):

you had the right idea except for the final step... once you found x= -1 or x=6 you find the y value (using either equation) and the (x,y) pairs are where the two curves intersect. i.e. are the "solutions"

OpenStudy (anonymous):

oh ok thank you i have to go :)